The weight of 72 books is 9 kg. What is the weight of 40 such books?

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Weight of 72 books = 9 kg
Weight of 1 book = 9 / 72
= 1 / 8 kg
So, weight of 40 books = (1 / 8) × 40 = 5 kg
∴Weight of 40 books is 5 kg

Thanks